Find the general solution of the following differential equation x dy / dt - 4y = x^5 e^x 


1
Expert's answer
2021-06-14T15:45:04-0400

"x\dfrac{dy}{dx}-4y=x^5e^x"

"y'-\dfrac{4}{x}y=x^4 e^x"

Find the integrating factor


"\mu(x)=\dfrac{1}{x^4}"

"(\dfrac{1}{x^4}y)'=e^x"

"\dfrac{1}{x^4}y=e^x+c_1"

"y=x^4e^x+c_1x^4"